OpenStudy (anonymous):

A particle is moving with velocity v(t) = t^2 – 9t + 18 with distance, s measured in meters, left or right of zero, and t measured in seconds, with t between 0 and 8 seconds inclusive. The position at time t = 0 sec is 1 meter right of zero, that is, s(0) = 1.

OpenStudy (anonymous):

The average velocity over the interval 0 to 8 seconds The instantaneous velocity and speed at time 5 secs The time interval(s) when the particle is moving right The time interval(s) when the particle is going faster slowing down Find the total distance the particle has traveled between 0 and 8 seconds

OpenStudy (anonymous):

For the first one I'd take the integral of the function from 0 to 8 right?

OpenStudy (anonymous):

@ganeshie8 Can you give this a look?

OpenStudy (anonymous):

well the instantaneous velocity at t=5 can be found by pluggin in t=5 the velocity function is the one given...

OpenStudy (anonymous):

also more math <.> Does this help? ^-^ a(t) = 2t - 9 ---> a(0) = -9 b/ x(t) = t^3/3 - 9t^2/2 + 18t + c , x(0)=1 ---> c=1 average velocity = (x(8)-x(0))/8 = ... c/ v(5) = 5^2 - 9(5)+18 = ...

OpenStudy (anonymous):

I see what you're doing but I don't think it's leading me to the right answer. You're saying the derivative of the velocity is equal to acceleration which is true. You also solved the integral but how does c=1? It is a definitive integral where C isn't included in the final answer no? I have a good amount of prior knowledge on the subject just need someone to point me to the right direction. Is there anything else you could share for the time intervals? That's the part that really makes me struggle.

OpenStudy (anonymous):

And thanks for actually helping, after 2 hours you'd think there'd be more responses.

OpenStudy (anonymous):

Yeah I'm sorry I'm Trying to explain slope.

OpenStudy (anonymous):

First one should be 224 According to @MaydayPaRAYde the second part should be equal to -2

OpenStudy (anonymous):

First on needs to be redone, just doesn't seem right to me.

OpenStudy (anonymous):

one*

OpenStudy (anonymous):

@IrishBoy123 do you have an idea?

OpenStudy (irishboy123):

this is lazy but at least it's the right way to go without a plotter, you should look at intercepts etc and plot it manually d is displacement naturally, usual caveat -- check the working for yourself then start answering......💥

OpenStudy (irishboy123):

https://www.desmos.com/calculator/hons6cwrqv i meant "this"

OpenStudy (anonymous):

Whoa..... hella confused now

OpenStudy (anonymous):

Basically graph and solve?

OpenStudy (anonymous):

Found my mistake for the first one. It'll be 26.66*1/8

OpenStudy (anonymous):

#1 = 3.33 #2 = v= -2 s=2

OpenStudy (irishboy123):

graphing really has to be part of it. this is applied maths. note however i did the integration as well BTW x on the graph represents time. i got funny stuff from desmos when i used t. i don't know desmos that well to have a quick fix.

OpenStudy (anonymous):

That's what I was planning to do, graph it to show the intervals and when it increase and decreases.

OpenStudy (anonymous):

I looked up how to do it on the web! I tried but honestly I'm not to this yet but I wanted to help you. I'm sorry that's how it taught me I am strugling o understand... :( I'm sorry

OpenStudy (anonymous):

It's cool, I am working through it. The points above the x-axis is where it is moving to the right, and the points below it is when it's too the left.

OpenStudy (anonymous):

Oh... That's Helpful.

OpenStudy (anonymous):

@IrishBoy123 what does the graph of the integral represent? displacement?

OpenStudy (irishboy123):

yes!!! displacement!! very important distinction

OpenStudy (anonymous):

Ok so I would use that for the last part right?

OpenStudy (anonymous):

Only problem now is the 4th one.

OpenStudy (anonymous):

Ohhhhh

OpenStudy (anonymous):

Derivative of velocity is equal to acceleration.

OpenStudy (irishboy123):

the distance is the area under the blue graph https://www.desmos.com/calculator/el3uwmejba there are a number of ways to get to it, i'm sure i don't know if you are integration but where as \(d(t) = \int dt \qquad v(t)\) \(|d(t)| = \int dt \qquad |v(t)|\) another reason why the plot is so useful.

OpenStudy (irishboy123):

sic *integration* integrating

OpenStudy (irishboy123):

soz by 4th you meant "The time interval(s) when the particle is going faster slowing down " yes use the graph or differentiate and solve

OpenStudy (irishboy123):

gtg for a while you seem in command

OpenStudy (anonymous):

Ight, just one more thing. Distance is always positive right?

OpenStudy (anonymous):

Lel Pretty sure it is.

OpenStudy (irishboy123):

yes, it is a scalar.

OpenStudy (irishboy123):

ie no direction, like the boy band.....😁

OpenStudy (anonymous):

Oh god.....that was terrible....

OpenStudy (anonymous):

Still having trouble or did you bump it on accident?

OpenStudy (anonymous):

Bumped for fun. I got this down.

OpenStudy (anonymous):

@IrishBoy123 Can't I just split the intervals then add? Like from 0 to 3, 3 to 6, 6 to 8?

OpenStudy (irishboy123):

yup

OpenStudy (anonymous):

That would give me distance correct?

OpenStudy (irishboy123):

yes

OpenStudy (irishboy123):

\[\text{Distance =} \int v + \int (-v) \text{[for when it's moving the other way]}+ \int v \qquad dt\]

OpenStudy (anonymous):

That's what I'm talking about! That's great!

OpenStudy (anonymous):

@IrishBoy123 35.667

OpenStudy (anonymous):

No need for you to check, I use my calculator for that one :p

OpenStudy (anonymous):

Fancy!

OpenStudy (irishboy123):

are we all agreed?!?! 🎃

OpenStudy (anonymous):

Yups only took 4hrs lel. Thanks for the help, really appreciate it. @IrishBoy123 @MaydayPaRAYde

OpenStudy (anonymous):

:) That's awesome I'm Glad he was here to put you on the right track!

OpenStudy (irishboy123):

thanks you mayday and ephemera 🍓🍓🍓

OpenStudy (anonymous):

You're Welcome :3